A study is conducted to investigate the impact of depression on a patient’s ability to survive cardiac disease. They evaluated 477 participants for depression and followed for 4 years. There were 374 patients with no depression (ND) and 69 of them died. Out of 103 patients with minor or major depression (D), 28 died.

Construct a 95% confidence interval for the difference in proportion of deaths between depressed patients with cardiac disease and non-depressed patients with cardiac disease. (Hint: Round sample proportions and SE to four decimal places.)

Question 6 options:

A)

(0.0392, 0.1356)

B)

(-0.0072, 0.1818)

C)

(0.0080, 0.1666)

D)

(-0.0371, 0.2118)

E)

(-0.0572, 0.2320)
Answer #1

ND 69 Pz=xz/n, = 0.1845 28 P:=xg/n = 0.2718 103 estimated difference in proportion 374 0.0873 = 1-Ø2 std error Se = V(.*(1-)/

option B is correct